OpenStudy (anonymous):

Choose the equation of the line passing through the point (-1, 7) and perpendicular to y = -negative one fourthx - 5. y = -4x - 11 y = -4x - 3 y = 4x + 11 y = 4x + 3

OpenStudy (amoodarya):

y = -negative one fourthx - 5 slope is -1/4=m when perpendicular m' *m=-1 so m'=+4

OpenStudy (amoodarya):

now write the line with slope +4 and passing through the point (-1, 7)

OpenStudy (anonymous):

according to my calculations, the answer is y=4x + 8 i used the slope-point formula

OpenStudy (anonymous):

so the slop is 4

OpenStudy (amoodarya):

y-y0=m'(x-x0) y-7=4(x-(-1)) y-7=4x+4 so y=4x+11

OpenStudy (anonymous):

thanks

OpenStudy (amoodarya):

your welcome